# Sea-Land-Cloud Segmentation in Satellite Hyperspectral Imagery by Deep Learning

### v1

@@ 2.3.1 Processing of 1D vs 3D Data
Our method involves training 16 distinct ML and DL models that are considered to be relevant for on-board processing, with a particular emphasis on DL models, as they are often better suited to capture complex data relationships. To train the models, we maintain the same data split used in our prior article [31], allocating 30 HS images for training (79%), 3 images for validation (8%), and 5 images for testing (13%). Furthermore, we employ an additional deployment set of 10 HS images, also from the sea-land-cloud dataset, to test the best-performing model’s ability to generalize to further new data. Additional details such as the specific image identifiers for the split are provided in the supplementary material referenced at the end of the paper.
@@
The ML and DL models presented in this work handle the data in either one-dimensional (1D) or three-dimensional (3D) formats. On the one hand, when we refer to 1D processing, it implies that the models make predictions solely based on spectral characteristics, disregarding any nearby spatial context. Examples of such models include 1D-CNNs based on 1D convolutions, where the input consists of a spectral signature without any spatial context. Henceforth, within this paper, we refer to these types of models as 1D models. When we refer to 3D processing, it signifies that the models concurrently analyze multiple adjacent spatial pixels, incorporating the hyperspectral datacube with both spectral (1D) and spatial (2D) information to predict the class for each pixel. Models like 2D-CNNs, that we refer henceforth as 2D models serve as examples, where the input is a 3D image containing both spatial and spectral data. Convolutional layers in 2D-CNN models operate by first applying 2D convolutions individually between each spectral channel and its corresponding kernel component. Subsequently, they aggregate the 2D convolutional results stacked in the spectral dimension to produce a final 2D feature map for the neural kernel.
Before introducing the models in the next section, we first describe the necessary arrangement of the data dimensions required for training them. On the one hand, in the case of 1D-CNN (or 1D ML) models, we transform each HS image into a 2D array resembling tabular data. Therefore, these models take inputs formatted as BATCH SIZE $\times$ CHANNELS, where BATCH SIZE indicates the number of image pixels for which predictions are to be inferred. 1D-CNN models perform 1D convolutions separately on each image pixel across its channels. On the other hand, for 2D-CNNs, optimizing memory and computational resources during training and inference is even more essential, and to help achieving this, we divide each HS image datacube into smaller cubes, often referred as 3D patches. First, to guarantee integer spatial dimensions for the patches, we explore both cropping and padding techniques prior to training. We eventually choose padding to retain the integrity of edge information in the images. Padding extends the image edges, which can be conveniently eliminated by cropping the padded spatial dimensions in the segmented image. After image padding, we proceed with non-overlapping patching. Unlike overlapping patching, this approach guarantees that each pixel is processed only once to reduce computational complexity. The input format for the 2D-CNNs is BATCH SIZE $\times$ PATCH SIZE $\times$ PATCH SIZE $\times$ CHANNELS, where BATCH SIZE refers in this case to the number of 3D patches. We find that a PATCH SIZE of 48 is suitable for our experiments, and this decision aligns with our design of the encoder-decoder networks, which consistently decrease the spatial dimensions by a factor of two via pooling. Consequently, selecting a PATCH SIZE of 48 ensures that the spatial dimensions subjected to pooling remain integer as the network’s depth increases.
@@ 2.3.2 Baseline Segmentation Models
As a starting point for on-board segmentation, we begin using simple baseline models consisting of traditional ML algorithms, often less complex and more explainable than deep neural networks. For convenience, during this initial phase, the models process 1D data (spectral signatures), which is a simpler processing strategy compared to handling 3D image patches. The initial baseline ML models consist of Stochastic Gradient Descent (referred henceforth as 1D-ML/SGD), Gaussian Naïve Bayes (1D-ML/NB), Linear Discriminant Analysis (1D-ML/LDA), and Quadratic Discriminant Analysis (1D-ML/QDA). Next, we incorporate multiple state-of-the-art DL models for 3D patch processing, such as NU-Net [35] and NU-Net-mod [12], previously employed in satellite imagery for cloud detection in RGB (FACSAT-1) and MSI (Sentinel-2B). Furthermore, we incorporate additional DL models, namely C-UNet and C-UNet++ [36], distinguished by the authors [36] as lightweight for their minimal parameter count and low computational requirements, making them well-suited for deployment on satellites. C-UNet and C-UNet++ are compact versions of the standard encoder-decoder U-Net architecture [37]. The subsequent literature [38] building upon [36], explores C-UNet and C-UNet++ models suggesting their possible use for segmentation in HS missions such as HYPSO-1. However, the method in [38] involves training on only two limited HS images, the conventional Pavia Centre and AeroRIT scene, which use sensors different from HYPSO-1 and do not include any cloud or oceanic data, which is of particular interest for ocean observation in HYPSO-1. In our work, we refer to the compact models in [38] as 2D-CUNet and 2D-CUNet++, and to their further reduced versions in [38] as 2D-CUNet Reduced and 2D-CUNet++ Reduced. Moreover, we add models from the FAUBAI project [39]. FAUBAI, a collaborative project between research and industry for ESA, focuses on segmentation for HS imagery for its possible future use in satellites like HYPSO-1, among other space missions. We refer to the networks in [39] as 2D-UNet FAUBAI and to its reduced version as 2D-UNet FAUBAI Reduced. Although Netteland [39] recommends any of these two networks for tasks such as cloud segmentation in HS satellite imagery, we note that similarly to [38], Netteland [39] conducted model testing only on a restricted set of four conventional HS images, namely, Pavia Centre, Pavia University, Salinas, and Indian Pines, which do not contain cloud or oceanic data.
2.3.3 Customised Segmentation Models: Lightweight CNNs for Optimal On-Board Deployment
2.3.3.1 Review of 1D-CNNs: Liuetal, Huetal, and LucasCNN Models
The Liuetal [40], Huetal [41] and LucasCNN [42] 1D-CNN models have been used before for data analysis in soil spectroscopy to segment various soil texture classes in HS data. We find these models of interest also for satellite-based inference due to their simple 1D processing approach. Indeed, in our previous article [31], we verified that a modified version of Liuetal achieved high performance, surpassing the state-of-the-art in soil spectroscopy [42].
@@ 2.3.3.1 Review of 1D-CNNs: Liuetal, Huetal, and LucasCNN Models
In Table I, we show the parameter counts (assuming 112 channels and 3 classes to segment) for the Liuetal, Huetal and LucasCNN models, highlighting that Liuetal has the lowest number of parameters among the three, although they still have parameter counts in the order of tens of thousands. Therefore, following our pursuit of further reducing the parameter count to merely a few thousands, more convenient for inference aboard satellites, we modify the soil models in [42], while aiming to preserve their performance, in the next sections.
@@
2.3.3.2 Hyperparameter Fine-Tuning Optimization
@@
To modify the soil models by reducing their trainable parameters while not degrading the models’ performance, we fine-tune their hyperparameters following the next procedure. To begin, in the first column of Table I, we list the state-of-the-art soil models in italic font [42], while our customized models are listed in bold. We also differentiate them by the addition of the term “Justo” to their names. In the subsequent columns of the table, for our models, we utilize bold font to highlight the hyperparameters that we modify compared to the soil models. Taking into account the promising nature of 1D-CNNs as lightweight models for on-board deployment, we focus the hyperparameter optimization exclusively on them, and we omit this optimization for 2D-CNN models. The modifications to hyperparameters, which are elaborated on in the following section, are a direct outcome of the subsequent hyperparameter fine-tuning process.
To optimize hyperparameters of 1D-CNNs, we begin by assigning them intuitive values that can yield acceptable performance. Then, we conduct a search using a hyperparameter grid, defined from the previous values, to identify the optimal combination of hyperparameters. This involves training and evaluating the resulting models with cross-validation and selecting the model with the highest overall accuracy. However, considering the high dimensionality of HSI data, we randomly choose only 2 million spectral signatures (approximately 10% of the training data), since training hundreds of models based on the multiple grid combinations would be exceedingly time-consuming, if the complete training set would have been used. Additionally, while evaluating the grid models through cross-validation, we restrict the evaluation to just two fold permutations to further decrease computation time. We note that it is not fully necessary to use the entire training set for hyperparameter tuning during grid search, as our goal is to obtain the optimal hyperparameters, yet within a reasonable time frame. Upon deciding the optimal hyperparameters, during the model training phase presented later, we certainly utilize all training pixels.
2.3.3.3 Introducing New Lightweigth 1D-CNNs: 1D-Justo-LiuNet, 1D-Justo-HuNet and 1D-Justo-LucasCNN Models
Figure 2: Lightweight 1D-Justo-LiuNet architecture for on-board semantic segmentation.
@@ 2.3.3.3 Introducing New Lightweigth 1D-CNNs: 1D-Justo-LiuNet, 1D-Justo-HuNet and 1D-Justo-LucasCNN Models
The hyperparameter optimization in previous section aims to achieve optimal accuracy, but we emphasize that the minimization of parameter counts is still the primary focus of our method. In this context, in our prior article [31], we presented a modified version of the Liuetal network resulting in 124,163 parameters, which might be considered large for in-flight deployment. However, our results demonstrated superior performance when contrasted with the Liuetal model [42]. We consider that the large parameter count in [31], which did not result in model overfitting, likely contributed to the significant performance improvement over the state-of-the-art. Nevertheless, in this work, our goal is to once again outperform the state-of-the-art, while also reducing the parameter count, in contrast to our previous work [31]. We aim to achieve a parameter count even lower than the count of the Liuetal model. Therefore, we next present the 1D-Justo-LiuNet (Fig. 2), which achieves 80% reduction in parameters compared to Liuetal (22,755) [42]. 1D-Justo-LiuNet is adapted to only 4,563 parameters, as shown in Table I. This reduction is accomplished by first following the approach in our prior article [31], suggesting to start with a base seed of neural kernels (32 in [31]) and progressively increase the kernel counts (denoted as N in Table I) based on the seed multiplied by factors of 2, 3, and 4 as the network deepens. In this work, we however start with a smaller seed of 6 kernels, instead of 32, aiming to achieve the substantial 80% reduction in the model’s parameter count. Moreover, as we did in our previous publication, we double the kernel size (denoted as K in Table I) compared to [42] in order to compensate for the reduction in the kernel counts, thereby ensuring that performance is not adversely affected.
@@
For the Huetal model [42], the high parameter count in Table I is mainly a result of a fully connected layer with numerous neurons preceding the output classification layer [41]. We refine the model, naming it 1D-Justo-HuNet, by reducing 70% of the neural units in this dense layer. By implementing additional modifications such as reducing the number of kernels and adjusting the kernel size, we attain a significant 90.50% reduction in parameter count (9,543). As for LucasCNN [42], our model, 1D-Justo-LucasCNN, follows the same approach applied in 1D-Justo-HuNet to reduce the parameter count, i.e., we significantly reduce the neural units in the two dense layers before the classification layer. Specifically, LucasCNN has 280 neural dense units, while 1D-Justo-LucasCNN reduces the units by 87% to just 35 neurons. Furthermore, 1D-Justo-LucasCNN includes other adjustments, such as transforming the 4-scale LucasCNN into a single-scale network, significantly reducing the number of kernels while tripling the kernel size, and employing a hyperbolic neural activation instead of a linear function. Table I illustrates a 68.41% reduction in parameter count (25,323) achieved by 1D-Justo-LucasCNN. The architectural details (as in Fig. 2) of the 1D-Justo-HuNet and 1D-Justo-LucasCNN are available in the supplementary materials, which also include the codes to reproduce the models.
2.3.3.4 Introducing a New Lightweight 2D-CNN: 2D-Justo-UNet-Simple Model
@@
Finally, to achieve lightweight models with reduced parameters, not only for 1D-CNN models as in previous section, but also for 2D-CNN architectures, we introduce: 2D-Justo-UNet-Simple (7,641 parameters in Table I). The 2D-Justo-UNet-Simple consists of a compressed version of the encoder-decoder U-Net structure to only 2 scales (as opposed to the typical 4 scales in U-Nets [37]), with a bottleneck level of only 12 kernels. Following each 2D convolution operation and before pooling and upsampling, we insert batch normalization before the rectified linear activation to enhance model generalization and mitigate potential gradient vanishing or exploding effects during training. We note that in the decoder’s contracting path, we employ upsampling as an alternative to utilizing operations like transposed convolutions. While transposed convolutions may enhance performance, they also lead to an increase in the number of trainable parameters. The code and architectural details (as in Fig. 2) of the 2D-Justo-UNet-Simple are provided in the supplementary materials.
@@ 2.3.4 Training Procedure
The DL models (implemented in Keras) are trained using an NVIDIA RTX A4000 GPU to increase the training speed, while the ML models (implemented in scikit-learn) are trained on a 13th Gen Intel(R) Core (TM) i9-13900K CPU, which results in a slower training process relying on CPU computation. We employ supervised learning to train the ML and DL models, with a particular focus on the training process for DL models. This is crucial because DL models are more susceptible to overfitting than ML models. We start by using the 112-channel dataset (from Section 2.2) to train the 1D and 2D models. However, when working with the 3-channel dataset, we exclusively focus on training 2D-CNN models. This decision is primarily based on the fact that 1D models do not take spatial information into account and rely solely on the spectral features, which, in this case, comprise only 3 channels. Furthermore, 1D-CNN models can face challenges related to the diminishing channel features as the network depth increases. When working with three channels, the reduction in features is so pronounced that, in the initial layers of the network, the processed features already diminish to zero, rendering prediction for three channels unfeasible.
@@
In our training procedure, we initially select high-level parameters like the number of epochs and the batch size, followed by the choice of the lower-level hyperparameters mentioned in previous sections. For both 1D-CNN and 2D-CNN models, we find that our dataset benefits from small batch sizes, increasing the frequency of parameter updates, but also extending the training time. In addition, we find that it is advisable to utilize a small number of epochs to prevent overfitting. Specifically, for 1D-CNNs, we use 2 epochs and a batch size of 32 pixels, while for 2D-CNNs, we opt for 3 epochs and a batch size of 4 patches. During training, we employ the Adam optimizer (Keras) with its default parameters to minimize the categorical cross-entropy loss, while using training accuracy as evaluation metric. As regards to batch size during inference, Keras dynamically selects the batch size for optimal memory efficiency.
@@
To evaluate the model’s performance at the end of each training epoch, we focus on the overall validation accuracy (validation set imbalance: 51.25%, 28.13%, and 20.62% for sea, land, and clouds, respectively), and we aim to achieve an empirical validation accuracy of around 0.90 while training. After completing the training process, we evaluate false positives and false negatives (inversely proportional) by treating the segmentation of each class as a binary classification problem. We note that the significance of these false positives and false negatives varies by class, as elaborated further in the upcoming section. However, we emphasize in advance that we aim for a reasonably balanced trade-off between false positives and false negatives across sea, land, and cloud classes before selecting the final model. In other words, we attempt to avoid situations where, for instance, we have very low false positives but extremely high false negatives. Therefore, we consider false positives and false negatives intuitively during model training, but we acknowledge more rigorous methods that focus on minimizing either false positives or false negatives by adjusting decision thresholds. These adjustments can involve, for instance, altering the decision boundary during the categorical class assignment based on different probability thresholds, consequently impacting the false positives and false negatives for each class. Additionally, applying penalties within the loss function to address specific misclassifications can also affect false positives and false negatives. However, in our work, we consider sufficient to predict the categorical classes during inference by selecting the class with the highest probability, while equally penalizing all classes during training, since the dataset imbalance is moderate. In the supplementary material, we provide the trained model files with the learned parameters in this work.
Finally, while quantization does contribute to further compacting lightweight models [11], we emphasize that we do not perform parameter quantization. The literature [14] demonstrates that in certain cases, where highly compressed light models are utilized, quantization is unnecessary, as the reduction in parameter count alone may be sufficient for in-orbit deployment.
2.4 Metrics for Segmentation Performance and Example Application: An On-Board Sea-Land-Cloud Ranking System
Figure 3: An intelligent on-board processing pipeline integrating a ranking system based on sea-land-cloud segmentation. $T\!H_{cl}$, $T\!H_{sea}$, and $T\!H_{land}$ represent thresholds configured at the operations center to regulate data management during in-flight operations.
To assess the models’ performance after training, we employ various metrics to evaluate the quality of the segmented results on the test set. First, our primary focus is on average accuracy, a widely used metric for evaluating segmentation models. Unlike overall accuracy, this metric is particularly convenient for imbalanced test sets, providing a more realistic assessment. Furthermore, adhering to the literature of cloud detection in satellite imagery, we also calculate the false and true alarm ratios. For this, as previously mentioned, we treat the segmentation of each class as a distinct binary classification problem. Finally, in our supplementary material further supporting the conclusions of our work, we present additional state-of-the-art metrics that we consider secondary in importance compared to the previously mentioned metrics. Specifically, these metrics include the F-1 score in addition to the overall accuracy.

### v2

@@ 2.3.1 Processing of 1D vs 3D Data
Our method involves training 16 distinct ML and DL models that are considered of interest for in-orbit deployment, with a particular emphasis on DL models, as they are often better suited to capture more complex data relationships. To train the models, we maintain the same data split used in our prior article [34], allocating 30 HS images for training (79%), 3 images for validation (8%), and 5 images for testing (13%). Furthermore, we employ a deployment set of 30 additional HS images, also from the sea-land-cloud dataset, to test the best-performing model’s ability to generalize to further new unseen data. Additional details such as the specific image identifiers for the data split are provided in the supplementary material (see data availability section).
@@
The ML and DL models presented in this work handle the data in either one-dimensional (1D) or three-dimensional (3D) formats. On the one hand, when we refer to 1D processing, it implies that the models make predictions solely based on spectral characteristics, disregarding any nearby spatial context. Examples of such models include 1D-CNNs based on 1D convolutions, where the input consists of a spectral signature with no spatial context. Henceforth, we refer to these types of models as 1D models. However, when we refer to 3D processing, it signifies that the models concurrently analyze multiple adjacent spatial pixels, incorporating the hyperspectral datacube with both spectral (1D) and spatial (2D) information to predict the class for each pixel. Models like 2D-CNNs, that we refer henceforth as 2D models serve as examples, where the input is a 3D image containing both spatial and spectral data. Convolutional layers in 2D-CNN models operate by first applying 2D convolutions individually between each spectral channel and its respective neural kernel component. Subsequently, they aggregate the 2D convolutional results stacked in the spectral dimension to produce a final 2D feature map. Finally, we describe in B the arrangement of the data dimensions required for training and inference. In the appendix, we offer e.g. information about how the images are divided into 3D patches for 2D-CNNs and the size of the patches used.
@@ 2.3.2 Baseline Segmentation Models
As a starting point, we begin using simple baseline models consisting of traditional ML algorithms as they are often more interpretable than deep neural networks. For convenience, during this initial phase, the models process 1D data (spectral signatures), which is a simpler processing strategy compared to handling 3D image patches. The initial ML baseline consists of Stochastic Gradient Descent (referred henceforth as 1D-ML/SGD), Gaussian Naïve Bayes (1D-ML/NB), Linear Discriminant Analysis (1D-ML/LDA), and Quadratic Discriminant Analysis (1D-ML/QDA). Next, we incorporate multiple state-of-the-art DL models for 3D patch processing, such as NU-Net [38] and NU-Net-mod [13], previously employed in satellite imagery for cloud detection in RGB (FACSAT-1) and MSI (Sentinel-2B). Furthermore, we incorporate additional DL models, namely C-UNet and C-UNet++, distinguished by the authors [39] as lightweight for their minimal parameter count and low computational requirements, making them well-suited for deployment on satellites. C-UNet and C-UNet++ are compact versions of the standard encoder-decoder U-Net architecture [40]. The subsequent literature [41] building upon [39], explores C-UNet and C-UNet++ models suggesting their possible use for segmentation in HS missions such as HYPSO-1. However, the method in [41] involves training on only two limited HS images, the conventional Pavia Centre and AeroRIT scene, which use sensors different from HYPSO-1 and do not include any cloud or oceanic data, which is of particular interest for ocean observation in HYPSO-1. In our work, we refer to the compact models in [41] as 2D-CUNet and 2D-CUNet++, and to their further reduced versions in [41] as 2D-CUNet Reduced and 2D-CUNet++ Reduced. Moreover, we add models from [42], which [42] refers to as the FAUBAI model with reference to the FAUBAI project that focuses on segmentation for HS imagery for its possible future use in satellites. We refer to the networks in [42] as 2D-UNet FAUBAI and to its reduced version as 2D-UNet FAUBAI Reduced. Although Netteland [42] recommends any of these two networks for tasks such as cloud segmentation in HS satellite imagery, we note that similarly to [41], Netteland [42] conducted model testing only on a restricted set of four conventional HS images, namely, Pavia Centre, Pavia University, Salinas, and Indian Pines, which do not contain cloud or oceanic data.
2.3.3 Customised Segmentation Models
2.3.3.1 Review of 1D-CNNs: Liuetal, Huetal, and LucasCNN
@@
The Liuetal [43], Huetal [44] and LucasCNN [45] 1D-CNN models have been used before for data analysis in soil spectroscopy to segment various soil texture classes in HS data. Additionally, in our previous article [34], we verified that a modified version of Liuetal achieved high performance in satellite HS imagery, surpassing the state-of-the-art in soil spectroscopy [45]. In this work, we find Liuetal, Huetal, and LucasCNN of interest also for in-orbit inference due to their simple 1D processing approach. Therefore, assuming 112 spectral channels and 3 classes to segment, in Table I we show the parameter counts for the three models. The parameter counts are in the order of tens of thousands with the Liuetal having the lowest number of parameters among the three. Following our pursuit of further reducing the parameter count to merely a few thousands, more convenient for inference aboard satellites, in the next sections we accordingly modify the soil models in [45], while aiming to preserve their performance.
@@
2.3.3.2 Ablation Study
@@
To adapt the soil models for deployment in space by reducing their trainable parameters while not degrading the models’ performance, we initially investigate the relevance of each neural layer and subsequently conduct hyperparameter fine-tuning. In the first column of Table I, we list the state-of-the-art soil models in italic font, while our customized models are listed in bold. We also differentiate our models by the addition of the term “Justo” to their names. In the subsequent columns of the table, for our models, we utilize bold font to highlight the hyperparameters that we modify compared to the soil models.
Taking into account the promising nature of 1D-CNNs as lightweight models for on-board deployment, we focus the hyperparameter optimization exclusively on them, and we omit the optimization for 2D-CNN models. The modifications to the hyperparameters in Table I further discussed in the following section, are a direct outcome of the hyperparameter fine-tuning process next. Namely, we begin by assigning them first intuitive values that can yield acceptable performance. Then, we conduct a search using a hyperparameter grid, based on the initial values, to identify the optimal combination of hyperparameters. To accommodate the high dimensionality of HSI data, we randomly choose only 2 million spectral signatures (approximately 10% of the training set) to reduce the time needed for the grid search. Additionally, while evaluating the grid models through cross-validation, we restrict the evaluation to only two fold permutations to decrease computation time further.
2.3.3.3 Introducing New Lightweigth 1D-CNNs:
1D-Justo-LiuNet, 1D-Justo-HuNet and 1D-Justo-LucasCNN
Figure 1: Lightweight 1D-Justo-LiuNet architecture for on-board segmentation. As an illustration, the block diagram employs 112 input channels for a single pixel. The network generates three detection probabilities, corresponding to sea, land, and clouds, respectively.
@@ 2.3.3.3 Introducing New Lightweigth 1D-CNNs:
The hyperparameter optimization process in previous section aims to achieve optimal accuracy, but we emphasize that the minimization of parameter counts for optimal in-orbit deployment is still the primary focus of our method. In this context, in our prior article [34], we presented a modified version of the Liuetal network resulting in 124,163 parameters, which might be considered large for in-flight deployment. However, our results demonstrated superior performance when contrasted with the Liuetal model [45]. We consider that the large parameter count in [34], which did not result in model overfitting, likely contributed to the significant performance improvement over the state-of-the-art. Nevertheless, in this article, our goal is to once again outperform the state-of-the-art, while also reducing the parameter count, in contrast to our previous work [34]. Indeed, we aim to achieve a parameter count even lower than the count of the Liuetal model in Table I. Therefore, we next present the 1D-Justo-LiuNet in Fig. 1, which achieves 80% reduction in parameters compared to Liuetal (22,755) [45]. 1D-Justo-LiuNet is adapted to only 4,563 parameters, as shown in Table I. This reduction is accomplished by first following the approach in our prior article [34], suggesting to start with a base seed of convolutional kernels and progressively increase the kernel counts (denoted as N in Table I) based on the seed multiplied by factors of 2, 3, and 4 as the network deepens. In this work, we however start with a small seed of 6 kernels, instead of 32 as in [34], to achieve the substantial 80% reduction in the model’s parameter count. Moreover, as we did in our previous publication, we double the kernel size (denoted as K in Table I) compared to [45] in order to compensate for the reduction in the kernel counts, thereby ensuring that performance is not adversely affected.
@@
For the Huetal model also in [45], the high parameter count in Table I is mainly a result of a fully connected layer with numerous neurons preceding the output classification layer. We refine the model, naming it 1D-Justo-HuNet, by reducing 70% the neural units in the fully connected layer. By implementing additional modifications such as reducing the number of convolutional kernels and adjusting their kernel size, we attain a significant 90.50% reduction in parameter count (9,543). As for LucasCNN [45], our model, 1D-Justo-LucasCNN, follows the same approach applied to Huetal to reduce the parameter count, i.e., we significantly reduce the neural units in the two dense layers before the classification layer. Specifically, LucasCNN has 280 neural dense units, while 1D-Justo-LucasCNN reduces the units by 87% to just 35 neurons. Furthermore, 1D-Justo-LucasCNN includes other adjustments, such as transforming the 4-scale LucasCNN into a single-scale network, significantly reducing the number of convolutional kernels while tripling the kernel size, and employing a hyperbolic neural activation instead of a linear function result of the hyperparameter optimization process. Table I illustrates a 68.41% reduction in parameter count (25,323) achieved by 1D-Justo-LucasCNN. The architectural details (as in Fig. 1) of the 1D-Justo-HuNet and 1D-Justo-LucasCNN are available in the supplementary materials, which also include the software codes to reproduce the models.
2.3.3.4 Introducing a New Lightweight 2D-CNN:
2D-Justo-UNet-Simple
@@ 2.3.3.4 Introducing a New Lightweight 2D-CNN:
Finally, to achieve lightweight models with reduced parameters, not only for 1D-CNN models as in previous section, but also for 2D-CNN architectures, we introduce: 2D-Justo-UNet-Simple with 7,641 parameters, as shown in Table I. The 2D-Justo-UNet-Simple consists of a compressed version of the encoder-decoder U-Net architecture with only 2 scales (as opposed to the typical 4 scales in U-Nets [40]), with a bottleneck level consisting of only 12 convolutional kernels. Additionally, following each 2D convolution operation and before pooling or upsampling, we insert batch normalization before the rectified linear activation to enhance model generalization and mitigate potential gradient vanishing or exploding effects during training. We note that in the decoder’s expanding path, we employ upsampling layers as an alternative to utilizing operations like transposed convolutions. While transposed convolutions may enhance performance, they also lead, however, to an increase in the number of trainable parameters. The software code and architectural details (as in Fig. 1) of the 2D-Justo-UNet-Simple are provided in the supplementary materials of this article.
@@ 2.3.4 Training Procedure
The DL models (implemented with Keras) are trained on an NVIDIA RTX A4000 GPU to increase the training speed, while the ML models (implemented with scikit-learn) are trained on a 13th Gen Intel(R) Core (TM) i9-13900K CPU, which results in a slower training process relying on CPU computation. We employ supervised learning to train the ML and DL models, with a special focus on the training process of DL models. This is crucial because, due to their larger number of parameters, DL models are more susceptible to overfitting than ML models.
@@
We start by using the 112-channel dataset (from Section 2.2) to train the 1D and 2D models. However, when working with the 3-channel dataset (from Section 2.2 and A), we exclusively focus on training 2D models instead of 1D models. This decision is primarily based on the fact that 1D models do not take spatial information into account and rely solely on the spectral features, which, in this case, comprise only 3 channels. Furthermore, 1D-CNN models can face challenges related to the diminishing channel features as the network depth increases. Namely, when working with only three channels, the reduction in features is so pronounced that, in the first layers of the network, the processed features already diminish to zero, rendering prediction for three channels unfeasible.
@@
In our training procedure, we initially select high-level parameters like the number of epochs and the batch size, followed by the choice of the lower-level hyperparameters previously discussed. In terms of high-level parameters, for both 1D-CNN and 2D-CNN models, we find that our dataset benefits from small batch sizes, increasing the frequency of updates in trainable parameters, but also extending the training time. As regards to batch size during inference, Keras dynamically selects the batch size for optimal memory efficiency. In addition, we find that it is advisable to utilize a small number of training epochs to prevent overfitting. Specifically, for 1D-CNNs, we use 2 epochs and a batch size of 32 pixels, while for 2D-CNNs, we opt for 3 epochs and a batch size of 4 patches. Furthermore, we employ the Adam optimizer with its default parameters to minimize the categorical cross-entropy loss, while using accuracy as evaluation metric in the training and validation sets.
@@
To evaluate the model’s performance at the end of each training epoch, we focus on the overall validation accuracy (validation set imbalance: 51.25%, 28.13%, and 20.62% for sea, land, and clouds, respectively), and we aim to achieve an empirical validation accuracy of around 0.90 during training. After completing the last training epoch, we evaluate false positives and false negatives (inversely proportional) by treating the segmentation of each class as a binary classification problem. In this context, we note that the significance of the false positives and false negatives varies per class, as elaborated further in the upcoming section. However, we note in advance that we aim for a reasonably balanced trade-off between false positives and false negatives across sea, land, and cloud classes before selecting the final trained model. Therefore, we consider false positives and false negatives intuitively during our training procedure, but we acknowledge more rigorous methods that focus on minimizing either false positives or false negatives by adjusting decision thresholds. The adjustments can involve, for instance, altering the decision boundary during the categorical class assignment based on different probability thresholds, consequently impacting the false positives and false negatives for each class. Additionally, applying penalties within the loss function to address specific class misclassifications can also affect false positives and false negatives. However, in our work, we consider sufficient to predict the categorical classes during inference by simply selecting the class with the highest probability, while equally penalizing all classes during training as the dataset imbalance is moderate. Finally, while quantization of trainable parameters certainly contributes to further compacting lightweight models [12], we emphasize that we do not perform parameter quantization in this article. The literature [15] demonstrates that in certain cases, where highly compressed light models are utilized (as in this work), quantization is unnecessary since the reduction in parameter count alone may be sufficient for in-orbit deployment. Furthermore, the effects of quantization can be highly dependent on the processing architecture.
In the supplementary material of this article, we provide the trained model files with the learned parameters (46 distinct files). Additionally, we include a user-friendly software program we develop, offering an intuitive graphical user interface for conducting oceanic-terrestrial-cloud segmentation on HYPSO-1 imagery.
2.4 Segmentation Performance Metrics and
Proposed Example Application
2.4.1 Proposed On-Board Sea-Land-Cloud Ranking System